OpenStudy (anonymous):

Identify the 8th term of the geometric sequence -2, -12, -72

OpenStudy (anonymous):

-3,359,232 -559,872 559,872 3,359,232 Answers

OpenStudy (anonymous):

-3,359,232

OpenStudy (anonymous):

r = common ratio = U2 / U1 = -12 / -2 = 6. a is the first term = -2 Un = a (r)^n-1 U8 = -2 (6)^8-1 = -2 (6)^7 =? * thanks to tcarroll for pointing that out